OpenStudy (anonymous):

What is 4x + 2y = -20 written in slope-intercept form? Answer y = 2x - 20 y = -4x + 20 y = 20x - 4 y = -2x - 10

OpenStudy (jhannybean):

To put this equation in slope intercept form,we will have to use the formula y = mx + b your equation is 4x + 2y = -20, subtract 4x from both sides to isolate y first

OpenStudy (anonymous):

Equation can be rewritten as 2y=-4x-20 Hence y=-2x-10 In General (u have to bring the equation in form of y=mx+c) m= slope & c= intercept here m=-2& c=-10

OpenStudy (jhannybean):

-4x + 4x + 2y = -20 - 4x Simplify 2y = -4x - 20 divide the whole equation by 2 2y/2 = -4x /2 - 20/2 now simplify y =-2x -10 This is your slope intercept form

OpenStudy (anonymous):

okay that makes sense thanks

OpenStudy (jhannybean):

No problem :)